OpenStudy (anonymous):

integrate (1-x) ln x dx

OpenStudy (turingtest):

by parts

OpenStudy (anonymous):

break it up into two integral (ln(x)) and -x(ln(x)) for the second part do a u substitution where u = ln(x); du = dx/x

OpenStudy (anonymous):

not sure if parts is necessary @TuringTest

OpenStudy (anonymous):

not using the udv = uv - vdu?

OpenStudy (anonymous):

I think parts is best here.

OpenStudy (turingtest):

yes, use dv=1-x and u=ln(x)

OpenStudy (anonymous):

can you show me the solution? i've done mine but it seems incorrect

OpenStudy (turingtest):

no we cannot give direct answers here, why don't you try and show us how you do?

OpenStudy (anonymous):

its pretty long. but my answer is ln x [ x - (x^2/2) ] - x - (x^2/4) + C

OpenStudy (anonymous):

yeah, u sub wouldn't work since u = ln(x), du = dx/x , x*du = dx.

OpenStudy (turingtest):

but that answer is correct :)

OpenStudy (turingtest):

oh I think except for a minus sign... ln x [ x - (x^2/2) ] - x + (x^2/4) + C

OpenStudy (anonymous):

thank you!!

OpenStudy (turingtest):

welcome :)